OpenStudy (bquiz):

t varies inversely with d and t = 0.15 when d = 120 {a} find t when d = 45 {b} find d when t = 0.12

OpenStudy (***[isuru]***):

\[t \ \ \alpha \ \ \frac{ 1 }{ d }\]

OpenStudy (***[isuru]***):

which means.. \[t = \frac{ k }{ d }\]

OpenStudy (***[isuru]***):

where k is a constant .. so u can write td = k which means that always the multiplication of t by d is a constant

OpenStudy (***[isuru]***):

since u know the value of t when d = 120 t= 0.15 and d = 120 u can use td = k and find the values of k

OpenStudy (***[isuru]***):

can u plug the values and find "k" ?

OpenStudy (***[isuru]***):

if u multiply 0.15 by 120 u will get 18.. so.. every time u multiply d by t the answer is 18 .. u can use this to find the answers for ur question

OpenStudy (***[isuru]***):

since the value u will get when u multiply any value for t or d by the corresponding value .. is 18 \[45t = 18\] and \[0.12 d = 18\] now solve these and u will get ur answer ... hope this will help you :)

OpenStudy (bquiz):

thanks

OpenStudy (***[isuru]***):

u r welcome :)